# Joint Demonstration and Preference Learning Improves Policy Alignment with Human Feedback

### v2

@@ 3 Proposed Approach: Alignment with Integrated Human Feedback (AIHF)
Figure 1: Comparison of the RLHF with the proposed AIHF.
3.1 Preliminaries
In this section, we develop our proposed formulation for the alignment problem. To flesh out our ideas, let us start with a brief discussion of the RLHF approach. First, the SFT stage maximizes the following likelihood function:
$\ell_{\rm SFT}(\pi):=\mathbb{E}_{\tau\sim\pi^{\rm E}}\Big{[}\log\prod_{t\geq 0}\Big{(}\pi(a_{t}|s_{t})\Big{)}^{\gamma^{t}}\Big{]}.$ (1)
For the RM stage, let $R(\tau;\theta):=\sum_{t\geq 0}\gamma^{t}r(s_{t},a_{t};\theta)$ denote the (parameterized) discounted cumulative reward of one trajectory $\tau$. Following the Bradley-Terry-Luce (BTL) model (Bradley and Terry, 1952), the human preference distribution for one pair of trajectories can be stipulated as (where $\sigma(\cdot)$ is the sigmoid function):
$P[\tau_{l}\prec\tau_{w}]=\sigma\big{(}R(\tau_{w};\theta)-R(\tau_{l};\theta)\big{)}.$ (2)
Based on (2), this stage maximizes the following:
$\ell_{\rm RM}(\theta):=\mathbb{E}_{(\tau_{l}\prec\tau_{w})\in\pi^{P}}\Big{[}\log\Big{(}\sigma\big{(}R(\tau_{w};\theta)-R(\tau_{l};\theta)\big{)}\Big{)}\Big{]}.$ (3)
Finally, the RL stage fine-tunes the SFT policy by solving the following policy optimization problem:
$\max_{\pi}\;\ell_{\rm RL}(\pi;R(\cdot;\theta)):=\mathbb{E}_{s_{0}\sim\rho,\tau\sim\pi}\bigg{[}R(\tau;\theta)+\sum_{t\geq 0}\gamma^{t}c(\pi)\bigg{]}$ (4)
where $c(\pi)$ is a penalty function, which can take the form of, e.g., the KL divergence $c(\pi):={D}_{\rm KL}\Big{(}\pi(\cdot|s_{t})\|\pi_{\rm SFT}(\cdot|s_{t})\Big{)}$. This penalty ensures the resulting policy is close to the original SFT policy. We have the following remark about the above problem.
Remark 3.1
(Optimal Policy for RLHF) Let $\pi_{\theta}$ denote the optimal policy for problem (4), where $\theta$ is used to parameterize the reward function. When $c(\pi):={D}_{\rm KL}\Big{(}\pi(\cdot|s_{t})\|\pi_{\rm SFT}(\cdot|s_{t})\Big{)}$, the KL-regularized MDP problem described by (4) has a closed-form solution:
$\pi_{\theta}(a|s)=\frac{\pi_{\rm SFT}(a|s)\exp(Q_{\theta}(s,a)/\beta)}{\sum_{\tilde{a}}\pi_{\rm SFT}(\tilde{a}|s)\exp(Q_{\theta}(s,a)/\beta)},$ (5)
$V_{\theta}(s):=\mathbb{E}_{\tau\sim\pi_{\theta}}\bigg{[}R(\tau;\theta)+\beta\sum_{t=0}^{\infty}\gamma^{t}{D}_{\rm KL}\Big{(}\pi(\cdot|s_{t})\|\pi_{\rm SFT}(\cdot|s_{t})\Big{)}\bigg{|}s_{0}=s\bigg{]}$ (6a)
Further, assuming that $T=1$, i.e., $\tau=(s_{0},a_{0})$, and considering the LLM alignment problem as a sequence-level training problem (this is a popular simplification in language models, see, e.g., Rafailov et al. (2023)), the closed-form expression of $\pi_{\theta}$ (5) can be reduced to:
$\pi_{\theta}(a_{t}|s_{t})=\frac{\pi_{\rm SFT}(a_{t}|s_{t})\exp(\frac{1}{\beta}r(a_{t},s_{t};\theta))}{\sum_{a_{t}\in A}\big{(}\pi_{\rm SFT}(a_{t}|s_{t})\exp(\frac{1}{\beta}r(a_{t},s_{t};\theta))\big{)}}.$ (7)
Remark 3.2
(Interpretation of the RLHF Policy) Let us take a closer look at (7). In this expression, the results of learning from demonstration data (i.e., $\pi_{\rm SFT}(\cdot)$) and from the preference data (i.e., $r(\cdot;\theta)$) are combined together in a product form to yield the final policy estimate. Such a form, though convenient, can result in a number of issues.
One such an issue is that, the reward $R(\tau;\theta)$ is learned only through the preference dataset $\mathcal{P}$. Despite the fact that problem (3) is relatively simple to optimize, such an approach has the following issues: 1) the high-quality demonstration data has been ignored in shaping the reward model; 2) in the situation where the preference data has limited coverage or is of low-quality, the learned reward model is likely to be of low quality as well. Here, low-quality preference data arises when, e.g., the trajectories are generated from low-quality models, or the preference annotation is noisy.
To see the impact on the noisy annotation, consider, for example, the case where a given $s_{t}$, suppose there are two actions $\hat{a}_{t}$ and $\tilde{a}_{t}$, and they are equally good. The SFT data respects the ground truth, and the learned SFT policy says $\pi_{\rm SFT}(\tilde{a}_{t}|s_{t})\approx\pi_{\rm SFT}(\hat{a}_{t}|s_{t})$. However, due to labeling noise, we have a preference data $(s_{t},\hat{a}_{t})\prec(s_{t},\tilde{a}_{t})$, therefore we have learned $R(a_{t},\tilde{a}_{t};\theta)>R(a_{t},\hat{a_{t}};\theta)$, yielding significant difference in the final policy.
@@ 3.2 A Meta-Formulation
The above discussion strongly motivates us to develop an approach that can model the entire alignment process with a common parametrization for both policy and reward models. Towards this end, let us consider the following meta-formulation:
${\bf(AIHF)}\quad\max_{\theta}$ $\quad L(\theta):=\alpha L_{1}(\pi_{\theta})+L_{2}(R(\cdot;\theta))$ (8a)
@@
where $\theta\in\mathbb{R}^{d}$ is a parameter; $L_{1}(\pi_{\theta})$ is a measure of fit of the parametrized policy $\pi_{\theta}$ to demonstration data and $L_{2}(R(\cdot;\theta))$ is a measure of fit of the parametrized reward model $R(\cdot;\theta)$ to the preference data and $L_{3}(\pi,R(\cdot;\theta))$ is a measure of performance of policy $\pi$ with respect to reward model $R(\cdot;\theta)$. and $\alpha\geq 0$ is one balancing coefficient reflecting the relative size of demonstration versus preference data. We refer the above formulation as Alignment with Integrated Human Feedback (AIHF); see Fig. 1 for an illustration of the proposed formulation. The AIHF (8) is a meta-problem that models the alignment problem. It has two levels: an upper-level problem in which the goal is to find policy and reward models that jointly maximize a measure of fit to demonstrations and preference datasets and a lower-level problem which ensures that the policy model optimizes performance with respect to the reward model. Its components can be customized to yield specific alignment formulations and algorithms. But before diving into various customizations, let us discuss the advantages of this formulation.
Generality. One can specialize the choice of loss functions and problem parameters, to yield a number of existing alignment formulations. Such generality implies that algorithms developed for (8) are easily applicable to different special formulations it covers. For more details see Sec. 3.4.
Joint optimization. The formulation jointly optimizes the reward and the policy. One benefit here is that it can strengthen the reward model through integrating both demonstrations and pairwise comparisons. Compared with the standard RLHF pipeline, through integrating additional data source such as demonstrations to train the reward model, it can further boost the policy optimization subroutine to achieve better alignment performance. See Sec. 5 for a detailed discussion on how the reward parameter $\theta$ is updated by leveraging such demonstration.
Dataset Integration. Clearly, the reward learning process leverages all the available data, therefore, we can expect that a high-quality reward model can still be obtained even under unfavorable situations where certain category of data is scarce or of low quality.
@@ 3.3 Specification of AIHF
$L_{1}(\pi_{\theta}):=\mathbb{E}_{\tau\sim\pi^{\rm E}}\Big{[}\sum_{t=0}^{\infty}\gamma^{t}\log\pi_{\theta}(a_{t}|s_{t})\Big{]}.$ (9)
Here $\pi_{\theta}$ optimizes the measure of performance $L_{3}(\pi;R(\cdot;\theta))$ for a reward model $R(\cdot;\theta)$ as below:
@@
$\mu_{\theta}:=\arg\max_{\mu\in\Delta^{T}}\mathbb{E}_{\tau\sim\mu}\big{[}R(\tau;\theta)-\beta\mathcal{D}_{KL}(\mu||\mu^{0})\big{]}$
@@
With this result, we can now obtain a model for the likelihood that sequence $\tau_{j}$ is preferred over $\tau_{j}$. By the independence of irrelevant alternatives property (Fudenberg et al., 2015) of the optimal choice $\mu_{\theta}$, when the set of feasible choices is reduced from $\mathcal{H}_{T}$ to just the the two-tuple $\{\tau_{l},\tau_{w}\}$, the likelihood that sequence $\tau_{w}$ is preferred over $\tau_{l}$ is given by $\mathbb{P}_{\theta}(\tau_{l}\prec\tau_{w}):=\frac{\mu_{\theta}(\tau_{w})}{\mu_{\theta}(\tau_{l})+\mu_{\theta}(\tau_{w})}.$ This motivates the choice of $L_{2}(\theta)$ as the following likelihood function:
$L_{2}(R(\cdot;\theta))=\mathbb{E}_{(\tau_{l}\prec\tau_{w})\in\pi^{P}}\Big{[}\log\frac{\mu_{\theta}(\tau_{w})}{\mu_{\theta}(\tau_{w})+\mu_{\theta}(\tau_{l})}\Big{]}=\mathbb{E}_{(\tau_{l}\prec\tau_{w})\in\pi^{P}}\Big{[}\log\frac{\mu^{0}(\tau_{w})\exp\big{(}R(\tau_{w};\theta)\big{)}}{\mu^{0}(\tau_{w})\exp\big{(}R(\tau_{w};\theta)\big{)}+\mu^{0}(\tau_{l})\exp\big{(}R(\tau_{l};\theta)\big{)}}\Big{]}.$
@@
${L}^{\rm BTL}_{2}(\theta)=\ell_{\rm RM}(\theta)=\mathbb{E}_{(\tau_{l}\prec\tau_{w})\in\pi^{P}}\Big{[}\log\Big{(}\sigma\big{(}R(\tau_{w};\theta)-R(\tau_{l};\theta)\big{)}\Big{)}\Big{]}.$ (11)
@@ 3.4 Special Cases of AIHF
Next we discuss how formulation (8) can be specialized to some of the known alignment algorithms.
Specialization to RLHF-Type Approach. First, if we set the coefficient $\alpha=0$ in (8), we obtain the following problem:
@@
Specialization to DPO-Type Approach. Let us consider the relationship between our formulation (8) with the DPO-type approaches. Let us set the following objective function $L_{1}=\ell_{\rm SFT}$ and $L_{2}=\ell_{\rm RM}$, and assume that $T=1$ for th generation process. Through relaxing the constraint which ensures the policy is optimal w.r.t. a certain parameterized model, we can obtain a DPO-type formulation:
$\max_{\pi}~{}L(\pi):=\alpha\cdot\mathbb{E}_{\tau^{\rm E}\sim\pi^{\rm E}}\bigg{[}\log\pi(a^{\rm E}|s^{\rm E})\bigg{]}+\mathbb{E}_{(\tau_{i}\prec\tau_{j})\sim\pi^{P}}\bigg{[}\sigma\big{(}\beta\log\frac{\pi(a_{j}|s_{j})}{\pi^{0}(a_{i}|s_{j})}-\beta\log\frac{\pi(a_{i}|s_{i})}{\pi^{0}(a_{j}|s_{i})}\big{)}\bigg{]}.$ (13)
@@
The above formulation specializes to Liu et al. (2024), which is a slightly generalized version of DPO when both demonstration and preference data are used. Setting $\alpha=0$ reduces to the problem solved by DPO; see Rafailov et al. (2023), Eq. (2).
4 Why AIHF can outperform two-stage alignment approaches?
To understand the difference between the proposed approach and the two-stage approach of the standard alignment pipeline, let us consider the following thought process over a simplified, but informative setting.
Let us consider the simplified case where there are $N$ discrete actions, that is, the action set is $A:=\{\tau_{1},\tau_{2},\cdots\tau_{N}\}$. Assume that there is a reward function $r(\cdot):A\mapsto\mathbb{R}$ (note, that here the parameterization of the reward has been omitted for simplicity). Let us compare the optimal solutions for different alignment algorithms and get some intuition about each approach. Throughout the process, we will assume that the reward function is not parameterized, and we directly learn the optimal reward value. Indeed this is a simplification needed to understand the structure of the solution, but we argue that this is without loss of generality, since if the reward model is sufficiently parameterized by neural networks, then these optimal reward values can always be recovered by proper training.
Consider the following decision-making model:
${\pi^{*}:=\arg\max_{\pi\in\Delta^{N}}\mathbb{E}_{\tau\sim\pi(\cdot)}\Big{[}r(\tau)-\beta\log\pi(\tau)\Big{]},}$ (14)
where $\Delta^{N}$ is the probability simplex and $\beta>0$. It can be shown that the solution to (14) is:
$\pi^{*}_{i}(r)=\frac{\exp(\frac{r_{i}}{\beta})}{\sum_{j}\exp(\frac{r_{j}}{\beta})},\quad\mbox{\rm where}\quad r_{i}:=r(\tau_{i}),\quad\forall~{}i.$ (15)
Reward Estimation via Demonstration Data: When demonstration data $\mathcal{D}:=\{\tau_{n}\}$ is available where each $\tau_{n}\in A,\;\forall~{}n$, the estimation problem can be written as:
$\widehat{r}_{\mathcal{D}}:=\arg\max_{r}\ell_{\mathcal{D}}(r)$
where $\ell_{\mathcal{D}}(r):=\mathbb{E}_{\tau_{i}\sim\mathcal{D}}[\log\pi^{*}_{i}(r)]$. By (15) it follows that
$\ell_{\mathcal{D}}(r)=\frac{1}{\beta}\mathbb{E}_{\tau_{i}\sim\mathcal{D}}\big{[}r_{i}\big{]}-\log\sum_{j}\exp(\frac{r_{j}}{\beta}).$
Hence, it is easy to verify the gradient is:
$\frac{\partial\ell_{\mathcal{D}}(r)}{\partial r_{i}}$ $=\frac{1}{\beta}\Big{(}\frac{\#\tau_{i}~{}\mbox{in}~{}\mathcal{D}}{|\mathcal{D}|}-\pi^{*}_{i}(r)\Big{)},~{}\forall~{}i.$ (16)
and the second order derivatives are:
$\frac{\partial^{2}\ell_{\mathcal{D}}(r)}{\partial r_{i}^{2}}=-\frac{1}{\beta^{2}}\pi^{*}_{i}(r)(1-\pi^{*}_{i}(r))$ $~{}~{}~{}~{}~{}\frac{\partial^{2}\ell_{\mathcal{D}}(r)}{\partial r_{i}\partial r_{j}}=\frac{1}{\beta^{2}}\pi^{*}_{i}(r)\pi^{*}_{j}(r).$ (17)
Therefore, we can observe that setting the gradient to zero, the first-order stationary solution (in fact, in this case it is optimal solution since the Hessian is negative definite) indicates that the estimate of the probability of choosing action $i$ is equal to the observed frequency of action $\tau_{i}$ appears in $\mathcal{D}$.
Note that in our simplified case, the above estimator can be viewed as the counterpart for the SFT estimator since it is the best that one can get using only demonstration data.
Reward Estimation with Preferences Data: Suppose that the preference data is available, $\mathcal{P}:=\{(\tau_{i}\prec\tau_{j})\}$, where again $\tau_{i},\tau_{j}\in A$. Let us understand how to analyze the optimal policy $\pi^{*}$. By IIA (independence of irrelevant alternatives) (Fudenberg et al., 2015), when the decision maker is asked to choose between only two alternatives $\{\tau_{i},\tau_{j}\}$, we have :
$\Pr(\tau_{j}\prec\tau_{i})$ $=\frac{\pi^{*}_{i}(r)}{\pi^{*}_{i}(r)+\pi^{*}_{j}(r)}.$
The reward estimation problem is defined as:
$\widehat{r}_{\mathcal{P}}\in\arg\max_{r}\ell_{\mathcal{P}}(r)$ (18)
where $\ell_{\mathcal{P}}(r):=\mathbb{E}_{(\tau_{j}\prec\tau_{i})\sim\mathcal{P}}\Big{[}\log\frac{\pi^{*}_{i}(r)}{\pi^{*}_{i}(r)+\pi^{*}_{j}(r)}\Big{]}$. The partial derivative of the log-likelihood is given by:
$\frac{\partial\ell_{\mathcal{P}}(r)}{\partial r_{i}}$ $=\frac{1}{\beta}\sum_{j\neq i}\big{(}\frac{\#\{\tau_{j}\prec\tau_{i}~{}\mbox{in}~{}\mathcal{P}\}}{|\mathcal{P}_{i,j}|}-\frac{\pi^{*}_{i}(r)}{\pi^{*}_{i}(r)+\pi^{*}_{j}(r)}\big{)}\frac{|\mathcal{P}_{i,j}|}{|\mathcal{P}|}.$ (19)
The first order condition can be written as:
$\pi^{*}_{i}(r)$ $=\frac{\sum_{j\neq i}\{\#a_{i}\succ a_{j}~{}\mbox{in}~{}\mathcal{P}\}}{\sum_{j\neq i}|\mathcal{P}_{i,j}|\rho^{*}_{-(i,j)}(r)}$ (20)

### v3

@@ 3 Alignment with Integrated Human Feedback (AIHF)
Figure 1: Comparison of the RLHF (left) with the proposed AIHF (right).
@@ 3.1 A Meta-Formulation
Towards developing an approach that can model the entire alignment process with a common parametrization for both policy and reward models, consider the following meta-formulation, termed Alignment with Integrated Human Feedback (AIHF):
${\bf(AIHF)}\quad\max_{\theta}$ $\quad L(\theta):=w_{1}L_{1}(\pi_{\theta})+L_{2}(R(\cdot;\theta))$ (5a)
@@
where $\theta\in\mathbb{R}^{d}$ is a parameter; $L_{1}(\pi_{\theta})$ is a measure of fit of the parameterized policy $\pi_{\theta}$ to demonstration data and $L_{2}(R(\cdot;\theta))$ is a measure of fit of the parameterized reward model $R(\cdot;\theta)$ to the preference data and $L_{3}(\pi,R(\cdot;\theta))$ is a measure of performance of policy $\pi$ with respect to reward model $R(\cdot;\theta)$. $w_{1}\geq 0$ is one balancing coefficient reflecting the relative size of demonstration versus preference data. Note that in the lower level policy optimization, the optimal policy corresponding to the reward model $R(\cdot;\theta)$ is actually determined by the reward parameter $\theta$, where we can denote the optimal policy under the reward model $R(\cdot;\theta)$ as $\pi^{*}_{R_{\theta}}:=\pi_{\theta}$. Therefore we simply put $\pi_{\theta}$ since the optimal policy under a certain reward model $R(\cdot;\theta)$ is actually determined by the reward parameters $\theta$.
See Fig. 1 for an illustration of AIHF. The AIHF (5) is a meta-problem that models the alignment problem. It has two levels: an upper-level problem in which the goal is to find policy and reward models that jointly maximize a measure of fit to demonstrations and preference datasets; and a lower-level problem which ensures that the policy model optimizes performance with respect to the reward model. Its components can be customized to yield specific alignment formulations and algorithms. Before diving into various customizations, let us discuss the advantages of this formulation.
Generality. One can specialize the loss functions and problem parameters to yield a number of existing alignment formulations. Such generality implies that algorithms developed for (5) are easily applicable to different special formulations it covers. For more details see Sec. 3.3.
Joint optimization. The formulation jointly optimizes the reward and the policy. One benefit here is that it can strengthen the reward model through integrating both demonstrations and pairwise comparisons. Compared with the standard RLHF pipeline, through integrating additional data source such as demonstrations to train the reward model, it can further boost the policy optimization subroutine to achieve better alignment performance. See Sec. 4 for a detailed discussion on how the reward parameter $\theta$ is updated by leveraging such demonstration, and see Tab. 5 for the experimental comparison between the reward model learned by RLHF and by our AIHF (5).
Dataset Integration. Clearly, the reward learning process leverages all the available data, therefore, we can expect that a high-quality reward model and its induced optimal policy can still be obtained even under unfavorable situations where the preference data is not sufficient. See Sec. 5 for experimental evidences.
@@ 3.2 Specification of AIHF
$L_{1}(\pi_{\theta}):=\mathbb{E}_{\tau\sim\mathcal{D}}\Big{[}\log\prod_{t\geq 0}\Big{(}\pi_{\theta}(a_{t}|s_{t})\Big{)}^{\gamma^{t}}\Big{]}=\mathbb{E}_{\tau\sim\mathcal{D}}\Big{[}\sum_{t\geq 0}\gamma^{t}\log\pi_{\theta}(a_{t}|s_{t})\Big{]}.$ (6)
Here $\pi_{\theta}$ optimizes the measure of performance $L_{3}(\pi;R(\cdot;\theta))$ for a reward model $R(\cdot;\theta)$ as:
@@
$\mu_{\theta}:=\arg\max_{\mu\in\Delta^{T}}\mathbb{E}_{\tau\sim\mu}\big{[}R(\tau;\theta)-\beta\mathcal{D}_{\rm KL}(\mu||\mu^{0})\big{]}$
@@
With this result, we can now obtain a model for the likelihood that sequence $\tau_{j}$ is preferred over $\tau_{j}$. By the independence of irrelevant alternatives property (Fudenberg et al., 2015) of the optimal choice $\mu_{\theta}$, when the set of feasible choices is reduced from $\mathcal{H}_{T}$ to just the the two-tuple $\{\tau_{l},\tau_{w}\}$, the likelihood that sequence $\tau_{w}$ is preferred over $\tau_{l}$ is given by $\mathbb{P}_{\theta}(\tau_{w}\succ\tau_{l}):=\frac{\mu_{\theta}(\tau_{w})}{\mu_{\theta}(\tau_{l})+\mu_{\theta}(\tau_{w})}.$ This motivates the choice of $L_{2}(\theta)$ as the following likelihood function:
$L_{2}(R(\cdot;\theta))$ $=\mathbb{E}_{(\tau_{w}\succ\tau_{l})\in\mathcal{P}}\Big{[}\log\frac{\mu_{\theta}(\tau_{w})}{\mu_{\theta}(\tau_{w})+\mu_{\theta}(\tau_{l})}\Big{]}$
$=\mathbb{E}_{(\tau_{w}\succ\tau_{l})\in\mathcal{P}}\Big{[}\log\frac{\mu^{0}(\tau_{w})\exp\big{(}R(\tau_{w};\theta)\big{)}}{\mu^{0}(\tau_{w})\exp\big{(}R(\tau_{w};\theta)\big{)}+\mu^{0}(\tau_{l})\exp\big{(}R(\tau_{l};\theta)\big{)}}\Big{]}.$
@@
${L}^{\rm BTL}_{2}(\theta)=\ell_{\rm RM}(\theta)=\mathbb{E}_{(\tau_{w}\succ\tau_{l})\in\mathcal{P}}\Big{[}\log\Big{(}\sigma\big{(}R(\tau_{w};\theta)-R(\tau_{l};\theta)\big{)}\Big{)}\Big{]}.$ (8)
@@ 3.3 Special Cases of AIHF
Next, we discuss how formulation (5) can be specialized to some of the known alignment algorithms.
Specialization to RLHF-Type Approach. First, if we set the coefficient $w_{1}=0$ in (5), we obtain:
@@
Specialization to DPO-Type Approach. Consider the relationship between formulation (5) with the DPO-type approaches. Let us set the following objective function $L_{1}=\ell_{\rm SFT}$ and $L_{2}=\ell_{\rm RM}$, and assume that $T=1$ for the generation process. Relaxing the constraint (5b) which ensures the policy is optimal w.r.t. a certain parameterized model, we can obtain a DPO-type formulation:
$\max_{\pi}~{}L(\pi):=w_{1}\cdot\mathbb{E}_{\tau^{\rm E}\sim\pi^{\rm E}}\bigg{[}\log\pi(a^{\rm E}|s^{\rm E})\bigg{]}+\mathbb{E}_{(\tau_{j}\succ\tau_{i})\sim\pi^{P}}\bigg{[}\log\bigg{(}\sigma\big{(}\beta\log\frac{\pi(a_{j}|s_{j})}{\pi^{0}(a_{i}|s_{j})}-\beta\log\frac{\pi(a_{i}|s_{i})}{\pi^{0}(a_{j}|s_{i})}\big{)}\bigg{)}\bigg{]}.$
@@
The above formulation specializes to Liu et al. (2024), which is a slightly generalized version of DPO when both demonstration and preference data are used. Setting $w_{1}=0$ reduces to the problem solved by DPO; see Rafailov et al. (2023, Eq. (2)), We refer to the algorithm defined by Eq. (5) as Direct AIHF.
Specialization to Self-Play Approach. Define $\ell(\cdot)$ as a monotonic and convex loss function, consider setting $L_{1}:=w_{1}\cdot\mathbb{E}_{\tau^{E}\in\pi^{E},\alpha\in\pi(.|s^{E})}\ell\left(R(\tau^{E};\theta)-R(\tau;\theta)\right)$, and setting $L_{2}$ and $L_{3}$ according to (8) and (7), respectively. Note that the choice of $L_{1}$ means that given demonstration data, we will find a policy that generates trajectories that match the rewards of the demonstration data. Again using the DPO type of reformulation, by substituting the reward expression obtained from the optimal policy (42) to $L_{1}$ and selecting the $\sigma(\cdot)$ as $\ell(\cdot)$, then the AIHF problem in this case becomes:
$\max_{\pi}L(\pi):=$ $w_{1}\mathbb{E}_{\tau^{\rm E}\sim\pi^{\mathrm{E}},\tilde{a}\sim\pi(\cdot|s)}\left[\log\sigma\left(\beta\log\frac{\pi(a^{\rm E}|s^{\rm E})}{\pi^{0}(a^{\rm E}|s^{\rm E})}-\beta\log\frac{\pi(\tilde{a}|s^{\rm E})}{\pi^{0}(\tilde{a}|s^{\rm E})}\right)\right]$
$+\mathbb{E}_{(\tau_{j}\succ\tau_{i})\sim\pi^{P}}\bigg{[}\log\sigma\left(\beta\log\frac{\pi(a_{j}|s_{j})}{\pi^{0}(a_{j}|s_{j})}-\beta\log\frac{\pi(a_{i}|s_{i})}{\pi^{0}(a_{i}|s_{i})}\right)\bigg{]}.$ (11)
Note that the first part of the above formulation is similar to what has been proposed in SPIN (Chen et al., 2024), which only utilizes the SFT data. We refer to the algorithm defined by Eq. (3.3) as Self-Play AIHF.
3.4 Why AIHF can outperform two-stage alignment approaches?
To understand the difference between the proposed approach and the successive stages approach of the standard alignment pipeline, let us consider the a static setting with action set is $A:=\{\tau_{1},\tau_{2},\cdots\tau_{N}\}$, reward function $R(\cdot):A\mapsto\mathbb{R}$, and demonstration $\mathcal{D}$ and preference dataset $\mathcal{P}$. In what follows, we compare the optimal solutions for policies obtained by different alignment approaches. Due to space limitation, all derivation in this section is relegated to Appendix A.2.
Policy with Demonstration Data. It can be easily shown that when only the demonstration data $\mathcal{D}$ is available, the probability of generating $i$-th data equals its empirical probability, i.e., $\pi_{\rm SFT}(\tau_{i})=\frac{\#\{\tau_{i}~{}\mbox{in}~{}\mathcal{D}\}}{|\mathcal{D}|}$. Assume that such a policy is parameterized by an implicit reward function $R_{D}$, using the following softmax choice model where $\tau_{i}\in A$ is selected with probability $\pi^{*}_{i}(R)=\frac{\exp(R_{i}/\beta)}{\sum_{j=1}^{N}\exp(R_{j}/\beta)}$ where $R_{i}:=R(\tau_{i})$. Assuming a reference value $\widehat{R}_{\mathcal{D}}(\tau_{1})=\bar{R}_{1}$, then the optimal rewards satisfies (See Sec. A.2.1):
#$\frac{\#\{\tau_{i}\in\mathcal{D}\}}{|\mathcal{D}|}=\pi^{*}_{i}(\widehat{R}_{\mathcal{D}})=\frac{\exp(\widehat{R}_{\mathcal{D}}(\tau_{i})/\beta)}{\sum_{j=1}^{N}\exp(\widehat{R}_{\mathcal{D}}(\tau_{j})/\beta)}~{}~{}~{}~{}i\in\{2,\dots,N\}.$ (12)
This implicit reward will be used shortly to characterize the RLHF policy.
Policy with Preference Only Data. Next, it can be shown that when only the preference data $\mathcal{P}$ is available, the reward estimation problem is defined as:
$\widehat{R}_{\mathcal{P}}=\arg\max_{R}\ell_{RM}(R):=\mathbb{E}_{(\tau_{i}\succ\tau_{j})\sim\mathcal{P}}\Big{[}\log\frac{\pi^{*}_{i}(R)}{\pi^{*}_{i}(R)+\pi^{*}_{j}(R)}\Big{]}.$ (13)
Again with a fixed reference value $\widehat{R}_{\mathcal{P}}(\tau_{1})=\bar{R}_{1}$, the solution is (see Sec. A.2.1):
$\pi^{*}_{i}(\widehat{R}_{\mathcal{P}})$ $=\frac{\sum_{j:j\neq i}|\mathcal{P}_{i\succ j}|}{\sum_{j:j\neq i}|\mathcal{P}_{i,j}|\rho_{-(i,j)}(\pi^{*}(\widehat{R}_{\mathcal{P}}))}$ (14)